Using the below information:

Allen Products, Inc. Income Statement for
the Year Ended December 31, 2019  
Sales revenue   $937,800
Less: cost of good sold   378,871
Gross profits   $558,929
Less: operating expenses   223,196
Operating profits   $335,733
Less: interest expense   32,823
Net profit before taxes   $302,910
Less: taxes (rate 30%)   90,873
Net profits after taxes   $212,037

Allen Products​ LP, wants to do a scenario analysis for the coming year. The pessimistic prediction for sales is $890,000​; the most likely amount of sales is $1,123,000​;

and the optimistic prediction is $1,271,000.

​Allen's income statement for the most recent year is shown above,

a. Use the ​percent-of-sales method, the income statement for December​ 31, 2019 and the sales revenue estimates to develop​ pessimistic, most​ likely, and optimistic pro forma income statements for the coming year.

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Sales:

Less: Cost of goods sold:

Gross profits:

Less: Operating expense:

Operating profits:

Less: Interest expense:

Net profits before taxes:

Taxes (30%):

Net profits after taxes:
Answer #1
Pessimatic most likely optimistic
Sales 890000 1123000 1271000
Less: cost of goods sold (40.40%) 359560 453692 513484
Gross profit 530440 669308 757516
Less: operating expenses (23.80%) 211820 267274 302498
Operating profits 318620 402034 455018
Less: interest expense (3.5%) 31150 39305 44485
Net profit before taxes 287470 362729 410533
Taxes (30%) 86241 108819 123160
Net profit after taxes 201229 253910 287373

Cost of goods sold = 378781/937800=40.40%

Operating expenses = 223196/937800=23.80%

Interest expense= 32823/937800=3.5%

Cost of goods sold = sales * 40.40%

Operating expenses = sales * 23.80%

Interest expense = sales *3.5%